# Beyond Memorization: Evaluating the True Type Inference Capabilities of LLMs for Java Code Snippets

### v1

@@ 3.1. StatType-SO
The StatType-SO dataset is prevalent for testing type inference tools (Phan et al., 2018; Saifullah et al., 2020; Dong et al., 2022; Huang et al., 2023b). It is constructed from real-world, manually repaired Java code snippets from Stack Overflow. StatType-SO contains 267 code snippets from six popular Java libraries (Android, GWT, Hibernate, JDK, JodaTime, XStream). The snippets were extracted from 50 randomly selected Stack Overflow posts for each library. To prepare the dataset, all parsing errors in the code snippets were first corrected. Next, the missing import statements were manually inferred and added to the code snippet, serving as the ground truth for evaluation.
@@
While StatType-SO is a widely adopted dataset for evaluating type inference techniques, it is important to acknowledge the potential for data leakage. Given its long-standing availability since 2017 at GitHub, StatType-SO may have been inadvertently incorporated into the training data of some LLMs, potentially leading to artificially inflated performance scores.
3.2. ThaliaType: A New Dataset
In this paper, we introduce ThaliaType, a new dataset specifically designed to rigorously evaluate the type inference capabilities of LLMs while mitigating the potential biases introduced by data leakage. ThaliaType is generated using Thalia, a tool originally developed for testing compilers’ type checkers. Recognizing its potential for LLM evaluation, we repurposed Thalia (Sotiropoulos et al., 2024) to generate a diverse set of syntactically valid, well-typed, and type-intensive Java programs. Given a set of types, along with the fields and methods defined for each type, Thalia generates programs that utilize a subset of these components. Each program consists of a single class with a main method containing variable declarations, field accesses, and method calls.
Thalia’s ability to generate syntactically correct and well-typed programs, combined with their type-intensive nature, aligns perfectly with the requirements for rigorous type inference evaluation. Thalia ensures that the programs are both valid and well-typed, eliminating the need for manual code cleanup or ground truth labeling. Furthermore, the type-intensive nature of these programs presents a meaningful challenge for type inference, while the inclusion of field accesses and method calls provides useful context for LLMs during inference.
@@ 3.2. ThaliaType: A New Dataset
To create ThaliaType, we used Thalia to generate 300 code snippets, with the types, fields, and methods drawn from the six libraries used in StatType-SO. For each library, 50 code snippets are generated. By selecting the same libraries, we ensured that the types in the generated code snippets were familiar to the LLMs, eliminating performance differences based on varying familiarity with the libraries. Figure 3 compares the statistics of ThaliaType and StatType-SO. The code snippets in both datasets are similar in terms of lines of code and the number of assignments. While Thalia-generated programs are type-intensive, ThaliaType includes fewer method calls but utilizes more types (via import statements) in each snippet.
@@
Our experiments show that SnR achieved comparable performance on both ThaliaType and StatType-SO (section 4.2), demonstrating that ThaliaType provides sufficient semantic information for the type inference task. To the best of our knowledge, ThaliaType is the first dataset designed for evaluating LLMs on code snippet type inference while addressing data leakage. Future research can leverage our replication package to generate additional code snippets using Thalia using their desired libraries to evaluate newer state-of-the-art LLMs.
@@
4. RQ1: How well do LLMs Perform Type Inference on Unseen Code Snippets
The data leakage problem is a real threat when evaluating LLM type inference performance, which motivated us to perform a thorough evaluation to uncover their true type inference capabilities. To mitigate this issue, in this RQ, we compared LLMs’ type inference performance on the publicly available StatType-SO dataset and the unseen code snippets in ThaliaType, providing insight into the potential impact of data leakage.
@@ 4.1. Method
Code snippets from StatType-SO and ThaliaType are given to LLMs using the prompt pattern in Figure 1 and also our baseline SnR one at a time without import statements. To check the output, import statements in LLMs’ response are extracted and compared against the original import statements in the code snippets. To ensure reproducibility, all the inferences are performed with a fixed seed (set to one) and a temperature of zero. GPT-4o-mini and GPT-4o are accessed using the OpenAI API (OPE, 2024b), while Llama3.1:8b and Llama3.1:70b are accessed through a self-hosted Ollama API (OLL, 2024) running on an A100 GPU.
The precision, recall, and F1-scores are computed to evaluate the performance of type inference. Precision measures the proportion of correctly inferred FQNs among all inferred FQNs, while recall measures the proportion of expected FQNs that are correctly inferred.
@@ 4.2. Results
LLMs achieved lower performance on unseen code snippets in ThaliaType despite their strong results in StatType-SO (Figure 4). Specifically, as shown in Figure 4(b), on the unseen ThaliaType code snippets, GPT-4o achieved a precision and recall of 54.74% and 44.54% respectively, a 42.8% and 53.1% decrease in precision and recall compared to results on StatType-SO. SnR outperformed all LLMs in ThaliaType with a precision of 84.15% and a recall of 84.43% . Even on StatType-SO, SnR achieved second best precision and recall of 95.50% and 91.46% . The large disparity between the LLMs’ performance on StatType-SO and unseen code snippets in ThaliaType suggests that LLMs are highly likely affected by data leakage which potentially inflated LLMs’ type inference performance on StatType-SO.
Unlike LLMs, SnR relies solely on the names of types, method names, and relationships between types in code snippets. The fact that LLMs outperformed SnR on StatType-SO suggests that LLMs leverage additional information not used by SnR to perform type inference. Conversely, the poor performance of LLMs on ThaliaType compared to SnR implies that LLMs failed to fully exploit the simple names, method calls, and type relations within the code snippet. These performance discrepancies motivated further investigations into the specific factors contributing to LLMs’ type inference capabilities. In particular, we explore in RQ2 (section 5) if LLMs leverage the available semantic information in code snippets or simply rely on matching exact syntax of the code snippets from training data.
@@ 4.2.1. Impact of Document Frequency on LLM Type Inference
To better understand why there is a large disparity between StatType-SO and ThaliaType, we investigated how the document frequency of a type impacts the performance of LLMs. Document frequency measures the number of source files that reference a specific type. Since LLMs’ outputs are biased by the training data (Huang et al., 2024; Liu et al., 2024), it is plausible that their performance may vary depending on the prevalence of a type within a LLMs’ training data. Understanding this bias helps ensure that LLMs deliver similar performance across all potential types used in the real world. To quantify a type’s document frequency, we analyzed source files from the Boa (Dyer et al., 2013a, b) GitHub dataset. Specifically, the document frequency is calculated as follows for a type .
@@
Based on the document frequency, we categorized the FQNs in the StatType-SO and ThaliaType datasets into distinct frequency ranges. For each group, we calculated the recall to investigate how a type’s real-world frequency impacts LLM performance. This categorization and analysis provide insights into the extent to which LLMs are biased toward more frequently encountered types.
@@
Results. LLMs performed poorly on less frequently used FQNs and worse on the unseen dataset ThaliaType than on StatType-SO, even for frequently used FQNs. To analyze how performance varies across FQNs with differing levels of usage frequencies, we grouped FQNs into frequency ranges by orders of magnitude (0-100, 100-1,000, 1,000-10,000, 10,000-100,000, and over 100,000 GitHub files). For the least frequently used FQNs in ThaliaType, GPT-4o only achieved a recall of 22.22% on StatType-SO and 10.36% on ThaliaType. In contrast, SnR demonstrated consistent performance regardless of a type’s popularity, achieving recalls of 100.00% and 73.94% on StatType-SO and ThaliaType respectively. On ThaliaType, LLMs only outperformed SnR on the most frequently used FQNs, specifically those appearing in over 100,000 GitHub source files. This disparity may arise from LLMs preferring more frequent FQN for a given simple name regardless of the context in the code snippet. For example, for a type with the simple name View, if android.view.View is the most frequent type in the ¿=1e5 category and javax.swing.text.View belongs to a less frequent category, then always recommending android.view.View will result in a high recall in the ¿=1e5 category but a lower recall in other categories.
@@
5. RQ2: To what extent do LLMs understand the semantics of code snippets?
@@
The performance decline on unseen data observed in RQ1 highlights the need to investigate how LLMs handle type inference. To explore whether this decline stems from a lack of semantic understanding in LLMs, we conducted an experiment using semantic-preserving transformations that alter syntax without changing the semantics. Our hypothesis is that if LLMs possess a deep understanding of the code snippets’ semantics, their performance should remain stable despite such transformations. Conversely, significant performance degradation would suggest that LLMs rely heavily on superficial syntactic patterns rather than a genuine semantic comprehension.
@@ 5.1. Method
The transformations, detailed in Algorithm 1, explore the extent to which syntactic changes affect LLMs’ semantic understanding of code. Three specific transformations were employed to assess the impact of modifications to identifier names (section 5.1.2), code structures (section 5.1.3), and Java keywords in comments (section 5.1.4). These were selected to highlight potential limitations in LLMs’ semantic understanding, though future research could explore additional transformations. Their impact was then evaluated using the methodology outlined in section 4. The Wilcoxon signed-rank test (Wilcoxon, 1992) was used to determine whether the transformations caused significant changes in precision, recall, and F1 scores.
@@ 5.1.1. Background: Java Grammar
The Java grammar depicted in Figure 5 is used to illustrate the key transformations on the Java code snippet. Each code snippet contains one or more classes along with a set of import statements. Classes may contain fields, methods, and blocks, with methods comprising blocks of statements and expressions. For simplicity, modifiers (e.g., public, static) were excluded, as they remain unchanged in the transformations. Other Java statements and expressions were similarly omitted as they follow the same transformation principles.
@@ 5.1.2. Identifier Renaming
This transformation investigates whether LLMs rely on specific identifier names for type inference. The RenameIdentifier function in Algorithm 1 details the process, which systematically renames identifiers such as variables, methods, classes, and packages in the code snippet. First, all variable declarations are identified using the FindVariableDeclaration function (Algorithm 1), which extracts statements matching the pattern . The algorithm then traverses each declaration, generating a unique, random three-word name for each variable using the GetRandomName function (Algorithm 1), from a predefined word list from prior work (Sotiropoulos et al., 2024) to ensure uniqueness. The algorithm replaces every occurrence of the variable name in the code (extracted using the getName function in Algorithm 1).
Next, the algorithm processes method declarations. To avoid renaming references that could alter semantics, all the method call expressions (s with the form ) that have the first optional part () is potentially external, thus added to a skip_list (Algorithm 1-1) to be filtered out (Algorithm 1). Additionally, overridden methods annotated with @Override are skipped, as their names must match those in the superclass (Algorithm 1). Lastly, the algorithm renames classes and packages (lines 1-1), following the same procedure used for variables. The details are omitted here for brevity.
@@
Figure 6 provides an example of this transformation, showing how variables, methods, classes, and packages are renamed. Method m which likely overrides the method m on type a is preserved out of an abundance of caution to preserve semantic consistency. Furthermore, it should be noted that different types of names are renamed separately. For example, the method name n is renamed to a different name than the variable name n.

### v2

@@ 3.1. StatType-SO
The StatType-SO benchmark suite is prevalent for testing type inference tools (Phan et al., 2018; Saifullah et al., 2020; Dong et al., 2022; Huang et al., 2023b). It is constructed from real-world, manually repaired Java code snippets from Stack Overflow. StatType-SO contains 267 code snippets from six popular Java libraries (Android, GWT, Hibernate, JDK, JodaTime, XStream). The snippets were extracted from 50 randomly selected Stack Overflow posts for each library. To prepare the benchmark suite, all parsing errors in the code snippets were first corrected. Next, the missing import statements were manually inferred and added to the code snippet, serving as the ground truth for evaluation. Figure 4 shows an example from StatType-SO. The code snippets from StatType-SO are generally short and make a small number of calls to the libraries and assign the intermediate results to variables.
@@
While StatType-SO is a widely adopted benchmark suite for evaluating type inference techniques, it is important to acknowledge the potential for data leakage. Given its long-standing availability since 2017 at GitHub, StatType-SO may have been inadvertently incorporated into the training data of some LLMs, potentially leading to artificially inflated performance scores.
3.2. ThaliaType: A New Benchmark Suite for Type Inference
We introduce ThaliaType, a new benchmark suite specifically designed to rigorously evaluate the type inference capabilities of LLMs while mitigating potential biases from data leakage. Code snippets in ThaliaType require reasoning over type-related semantic information to accurately infer the types used. Through this design, ThaliaType evaluates how effectively LLMs can leverage type information in code snippets, ensuring that the measured performance is generalizable to previously unseen code. If an LLM understands the type information provided, then it can correctly infer the types for any given code snippet.
ThaliaType is generated using Thalia (Chaliasos et al., 2022; Sotiropoulos et al., 2024), a tool originally developed for testing the type system of compilers. Thalia has uncovered 117 confirmed or fixed type-related bugs across four major compilers. This large number of bugs demonstrates that compiler developers value those found bugs though triggered by automatically generated programs, which indirectly demonstrates the typing relations in Thalia-generated code snippets might resemble the typing relations in real-world code snippets. Recognizing its value for LLM evaluation, we repurposed Thalia to create a diverse set of syntactically valid, well-typed Java programs. Given a set of types, fields, and methods, Thalia systematically generates programs that utilize a subset of these components. By reasoning with the type information within the code snippets, LLMs can infer the types used in ThaliaType.
The general process for generating ThaliaType code snippets is summarized in Algorithm 1. Six libraries that were used in StatType-SO were given to Thalia as input. By selecting the same libraries, we ensured that the types in the generated code snippets were familiar to the LLMs, eliminating performance differences caused by varying familiarity with the libraries. For each library, 50 code snippets are generated. Each snippet is based on a single library, following the convention in StatType-SO. In total, 300 code snippets were generated.
Figure 5 shows an example code snippet generated by ThaliaType. Each code snippet consists of a Main class and a test method. In this code snippet, the code creates a new LocalTime from a given time represented by a long value of -58. The chronology from this LocalTime, which in JodaTime is ISOChronology, is then used to construct the current DateMidnight. Finally, the getDayOfYear() method is called to retrieve the ordinal day number of the year.
@@ 3.2. ThaliaType: A New Benchmark Suite for Type Inference
Similarity between ThaliaType and StatType-SO. The code snippets in ThaliaType are similar to the code snippets from StatType-SO. Figure 6 compares the statistics of the two benchmark suites. The code snippets in both suites are similar in terms of lines of code and the number of assignments. While Thalia-generated programs are type-intensive, ThaliaType includes fewer method calls but utilizes more types (via import statements) in each snippet. More importantly, our experiments show that SnR achieved comparable performance on both ThaliaType and StatType-SO (section 5.2), demonstrating that ThaliaType provides as much semantic information as StatType-SO for the type inference task. If LLMs genuinely understand the semantics of code snippets, LLMs should be able to perform similarly on both suites.
@@
To the best of our knowledge, ThaliaType is the first benchmark suite designed for evaluating LLMs on code snippet type inference while addressing data leakage. Future research can leverage our replication package to generate additional code snippets using Thalia using their desired libraries to evaluate newer state-of-the-art LLMs.
4. Data Leakage in LLM-Based Type Inference
In this section, we aim to investigate whether the evaluation of LLM-based type inference techniques may be affected by the data leakage issue (Inan et al., 2021). Data leakage occurs when the training datasets overlap with the benchmark datasets used for evaluation, leading models to achieve artificially high performance by memorizing examples rather than demonstrating genuine generalization. The evaluation of the prior studies that either compared against (Huang et al., 2023a; Chen et al., 2024) or incorporated LLMs (Kabir et al., 2025) largely relied on the StatType-SO dataset as the benchmark suite, which has been publicly available on GitHub since 2017 (Sta, 2017). Because the GPT-based models used in these evaluations were likely trained on datasets containing StatType-SO, their reported performance may be affected by data leakage.
However, directly detecting data leakage from LLMs’ training is challenging, as most datasets are proprietary and undisclosed, with only a few released to the broader community. Two widely used LLM families, GPT (OpenAI, 2024) and Llama (Meta, 2024), are trained on proprietary datasets with only partial disclosure of their training data. GPT’s creator, OpenAI, has stated that the models are trained on a mixture of publicly available data (such as internet text) and data licensed from third-party providers. However, the exact composition of the dataset remains undisclosed, and access to the model is limited to APIs. Similarly, while Llama models are open-weight and thus freely available for download and use, the data used to train them is not publicly released. Meta discloses that Llama is trained on publicly available data, primarily sourced from the internet, but the exact sources are not specified. This lack of transparency makes it difficult to assess the risk of data leakage in either model.
In contrast, StarCoder2 (Lozhkov et al., 2024) is trained entirely on the openly available dataset StackV2 (BigCode, 2024), which comprises 67.5TB of source code across 658 programming languages. Because this dataset is openly accessible, it is possible to directly examine the training corpus and assess the extent of any data leakage.
| Repository Owner | Repository Name | # of code snippets | Last Commit Date |
| miketran238 | AndroidOracle | 50 | 2017-02-14 |
| miketran238 | JodatimeOracle | 50 | 2017-02-12 |
| mrthlinh | Oracle-GWT | 50 | 2017-02-13 |
| mrthlinh | Oracle-Hibernate | 50 | 2017-02-13 |
| mrthlinh | TypeResolution_Oracle | 219 | 2017-02-12 |
| pdhung3012 | TypeResolution_Oracle | 267 | 2017-06-30 |
To investigate the content of the StackV2, the metadata of the files are examined. Using the matching procedure described in Algorithm 2, file names in StackV2 were compared against those in the StatType-SO benchmark suite, which identified repositories containing overlapping code snippets. This process produced 26 repositories and three repository owners with five or more matching file names, forming the basis for manual inspection.
Manual analysis identified six repositories (shown in Table 2) containing code snippets alongside their ground truth from the StatType-SO benchmark suite. Notably, the TypeResolution_Oracle repository by mrthlinh includes 219 code snippets with ground truth, excluding only those dependent on the Android library. In addition, pdhung3012’s TypeResolution_Oracle repository contains the complete StatType-SO benchmark suite. All identified repositories were committed in 2017, well before the knowledge cutoffs of GPT and Llama, making it highly likely that these models used for evaluation were trained on the full StatType-SO benchmark suite.
@@ 4. Data Leakage in LLM-Based Type Inference
5. RQ1: How well do LLMs perform type inference on unseen code snippets?
Given the strong possibility of data leakage, we sought to rigorously assess the genuine type inference capabilities of LLMs. To do this, RQ1 investigates their performance on two benchmark suites: the original StatType-SO (which may have been seen during training), and ThaliaType (which contains only unseen code snippets). This comparison allows us to isolate the impact of potential training overlap and better understand how well these models generalize to genuinely novel code.
@@ 5.1. Method
Code snippets from StatType-SO and ThaliaType are given to LLMs using the prompt pattern in Figure 3 and also our baseline SnR, one at a time, without import statements. To check the output, import statements in LLMs’ response are extracted and compared against the original import statements in the code snippets. To ensure reproducibility, all the inferences are performed with a fixed seed (set to one) and a temperature of zero. GPT-4o-mini and GPT-4o are accessed via the OpenAI API (OPE, 2024b), while StarCoder2:15b, Llama3.1:8b, and Llama3.1:70b are accessed through a self-hosted Ollama API (OLL, 2024) running on an A100 GPU. The total cost of using the GPT models for RQ1 was $4.47, based on OpenAI’s pricing as of early 2025.
Precision, recall, and F1-scores are computed to evaluate the performance of type inference. Precision measures the proportion of correctly inferred FQNs among all inferred FQNs, while recall measures the proportion of expected FQNs that are correctly inferred.
@@ 5.2. Results
Despite achieving strong results in StatType-SO (Figure 7), all evaluated LLMs demonstrated lower performance on unseen code snippets in ThaliaType. StarCoder2:15b consistently ranked between Llama3.1:8b and Llama3.1:70b across both benchmark suites. Specifically, as shown in Figure 7(a), StarCoder2:15b achieved an F1 score of 81.67% on StatType-SO, higher than Llama3.1:8b, but lower than Llama3.1:70b, GPT-4o, and GPT-4o-mini. However, its performance dropped sharply to 27.08% on ThaliaType, representing a 66.8% decrease. Despite this decline, it still outperformed Llama3.1:8b but remained below the other larger models. Notably, even the best performing model, GPT-4o, experienced a 48.5% decrease in F1 score when evaluated on ThaliaType. The consistent performance drop across all models, including StarCoder2:15b with confirmed data leakage, suggests that LLMs are likely affected by data leakage, which potentially inflated LLMs’ type inference performance on StatType-SO.
Unlike LLMs, SnR relies solely on the names of types, method names, and relationships between types in code snippets. SnR only experienced a 9.8% decrease in F1 score when evaluated on ThaliaType, due to the difference in available information in the code snippets. The fact that LLMs outperformed SnR on StatType-SO while performing poorly on ThaliaType suggests that LLMs leverage additional information in StatType-SO to perform type inference. These performance discrepancies motivated further investigations into the specific factors contributing to LLMs’ type inference capabilities. In particular, in RQ2 (section 6), we investigate what types of information LLMs might leverage for type inference, and whether their performance depends on matching exact syntax of the code snippets seen during training.
@@ 5.2.1. Impact of Document Frequency on LLM Type Inference
To better understand the large performance disparity between StatType-SO and ThaliaType, we investigated how the document frequency of a type impacts the performance of LLMs. Document frequency measures the number of source files that reference a specific type. Since LLM outputs are biased by training data (Huang et al., 2024; Liu et al., 2024), their performance may vary depending on the prevalence of a type in the training corpus. Understanding this bias helps ensure that LLMs deliver similar performance across all potential types used in the real world. To quantify a type’s document frequency, we analyzed source files from the Boa (Dyer et al., 2013a, b) GitHub dataset. Specifically, the document frequency is calculated as follows for a type .
@@
Based on the document frequency, we categorized the FQNs in the StatType-SO and ThaliaType benchmark suites into distinct frequency ranges. For each group, we calculated the recall to investigate how a type’s real-world frequency impacts LLM performance. This categorization and analysis provide insights into the extent to which LLMs are biased toward more frequently encountered types.
@@
Results. LLMs performed poorly on less frequently used FQNs and worse on the unseen benchmark ThaliaType than on StatType-SO, even for frequently used FQNs. To analyze how performance varies across FQNs with differing usage frequencies, we grouped FQNs into frequency ranges by orders of magnitude (0-100, 100-1,000, 1,000-10,000, 10,000-100,000, and over 100,000 GitHub files). The results, shown in Table 3, indicate that LLM performance degrades as the document frequency of FQNs decreases. Specifically, for the least frequently used FQNs, GPT-4o only achieved a recall of 22.22% on StatType-SO and 10.36% on ThaliaType. In contrast, SnR demonstrated consistent performance regardless of a type’s popularity, achieving recalls of 100.00% and 73.94% on StatType-SO and ThaliaType respectively. On ThaliaType, LLMs only outperformed SnR on the most frequently used FQNs, specifically those appearing in over 100,000 GitHub source files. This disparity may arise from LLMs preferring more frequent FQNs for a given simple name regardless of the context in the code snippet. For example, for a type with the simple name View, if android.view.View is the most frequent type in the ¿=1e5 category and javax.swing.text.View belongs to a less frequent category, then always recommending android.view.View will result in a high recall in the ¿=1e5 category but a lower recall in other categories.
Interestingly, despite StarCoder2:15b being trained on code snippets from StatType-SO, it achieved a maximum recall of only 87.26% . As in the previous analysis, its performance generally fell between Llama3.1:8b and Llama3.1:70b on both StatType-SO and ThaliaType. Although training data may inflate performance, StarCoder2:15b does not reproduce code snippets verbatim, as models are designed to generalize beyond specific examples. This generalization, while desirable, makes detecting instances of data leakage difficult.
@@
6. RQ2: To what extent do LLMs understand the execution semantics of code snippets?
@@
The performance decline on unseen data observed in RQ1 highlights the need to investigate how LLMs perform type inference. To explore whether this decline stems from a lack of semantic understanding in LLMs, we conducted an experiment using semantic-preserving transformations that alter syntax without changing the semantics. Our hypothesis is that if LLMs possess a deep understanding of code snippet semantics, their performance should remain stable despite such transformations. Conversely, significant performance degradation would suggest that LLMs rely heavily on superficial syntactic patterns rather than a genuine semantic comprehension.
@@ 6.1. Method
The transformations, detailed in Algorithm 3, explore the extent to which syntactic changes affect LLMs’ semantic understanding of code. Three specific transformations were employed to assess the impact of modifications to identifier names (section 6.1.2), code structures (section 6.1.3), and Java keywords in comments (section 6.1.4). These were selected to highlight potential limitations in LLMs’ semantic understanding, though future research could explore additional transformations. Their impact was then evaluated using the methodology outlined in section 5. The Wilcoxon signed-rank test (Wilcoxon, 1992) was used to determine whether the transformations caused significant changes in precision, recall, and F1 scores. The average running cost per transformation for GPT models was $6.30 in RQ2, slightly higher than in RQ1 because of longer code snippets.
@@ 6.1.1. Background: Java Grammar
The Java grammar depicted in Figure 8 is used to illustrate the key transformations on the Java code snippet. Each code snippet contains one or more classes along with a set of import statements. Classes may contain fields, methods, and blocks, with methods comprising blocks of statements and expressions. Highly repetitive Java statements and expressions were omitted as they follow the same transformation principles.
@@ 6.1.2. Identifier Renaming
This transformation investigates whether LLMs rely on specific identifier names for type inference. Although identifier names can provide useful signals, overreliance on unique identifiers or specific identifier sequences seen during training may limit a model’s ability to generalize. In real-world code snippets, identifiers are often lowercase variants of type names or abbreviated forms, providing no additional semantic information.
The RenameIdentifier function in Algorithm 3 details the process, which systematically renames identifiers such as variables, methods, classes, and packages in the code snippet. First, all variable declarations are identified using the FindVariableDeclaration function (Algorithm 3), which extracts statements matching the pattern . The algorithm then traverses each declaration, generating a unique, random three-word name for each variable using the GetRandomName function (Algorithm 3), from a predefined word list from prior work (Sotiropoulos et al., 2024) to ensure uniqueness. The algorithm replaces every occurrence of the variable name in the code (extracted using the getName function in Algorithm 3).
Next, the algorithm processes method declarations. To avoid renaming references that could alter semantics, all the method call expressions (s with the form ) that have the first optional part () are potentially external, thus added to a skip_list (lines 3-3) to be filtered out (Algorithm 3). Additionally, overridden methods annotated with @Override are skipped, as their names must match those in the super class (Algorithm 3). Lastly, the algorithm renames classes and packages (lines 3-3), following the same procedure used for variables. The details are omitted here for brevity.
@@
Figure 9 provides an example of this transformation, showing how variables, methods, classes, and packages are renamed. The method m declared at line 5 likely overrides a method on the super class, A.m(). Out of an abundance of caution to preserve semantic consistency, m is not renamed. Furthermore, it should be noted that different types of names are renamed separately. For example, the method name n is renamed differently from the variable name n.